OpenStudy (anonymous):

How long will it take for $1700 to grow to $38,500 at an interest rate of 8.8% if the interest is compounded continuously? Round the number of years to the nearest hundredth?

OpenStudy (amistre64):

38500 = 1700(e^(.088t))

OpenStudy (amistre64):

385/17 = e^.088t

OpenStudy (amistre64):

ln(385/17) = .088t

OpenStudy (amistre64):

ln(385/17)/.088 = t

OpenStudy (amistre64):

ln(385 / 17) / .088 = 35.4548863 is what google gives me

OpenStudy (anonymous):

A = P*e^(rt) 38,500 = 1700e^(0.088t) (38,500/1,700) = e^(0.088t) 0.088t = ln(38,500/1,700) t = (1/0.088)*ln(38,500/1,700) t≈35.45